OpenStudy (anonymous):

find the y-intercept(s),if any for the equation:y=x2-16?

OpenStudy (anonymous):

just plug x =0 there to get y, that is y-intercept

OpenStudy (eric_d):

y=mX+b b=y-intercept So,it's -16

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

it's \(y= x^2 -16\) , it is a parabola, not a line with the standard equation of y = mx +b